OpenStudy (anonymous):

what is the value of the derivative of y=(x^5)(lnx) when x=4

OpenStudy (loser66):

what do you get so far?

OpenStudy (anonymous):

f'(x)= x^(5)= n( X^(n-1))= 5X^(4)= 5(4)^4)(1/4)

OpenStudy (anonymous):

f(x)= ln(x) is (1/x)

OpenStudy (loser66):

waaaattt?

OpenStudy (loser66):

is it not a product rule? how can you separate it into 2 parts like this?

OpenStudy (anonymous):

wow! Your right! Sorry lead you the wrong way!

OpenStudy (anonymous):

\[f(g(x))'=f'(g(x)g'(x), f(x)=x^5 and g(x)= \ln( x) \]

OpenStudy (anonymous):

right ?

OpenStudy (loser66):

nope,

OpenStudy (loser66):

it's not f(g(x)) , it's f(x)* g(x)

OpenStudy (anonymous):

\[my calculator says = (5\ln(l/x)-1 ) (x^4)\]

OpenStudy (loser66):

my brain says the same answer with you

OpenStudy (anonymous):

confusion! hahaha product rule right ? i"m getting (5x^4)(lnx)+(x^5)/x